# Nash CoT: Multi-Path Inference with Preference Equilibrium

### v1

@@ Abstract
Chain-of-thought (CoT) prompting has emerged as a powerful technique for enhancing the reasoning capabilities of Large Language Models (LLMs) on complex problems. Among CoT-related studies, self-consistency (Multi-path inference with answer filtering through voting) involves generating multiple reasoning paths using the CoT framework and then selecting the most frequently produced outputs standing out as a concise yet competitive approach. While self-consistency has indeed led to the improvements in LLM inference, the use of multi-path inference also escalates deployment costs. Therefore, maintaining the performance benefits of self-consistency inherited from multi-path inference while reducing the inference costs holds significant value. In this research, we conceptualize language decoding as a preference consensus game, constructing a bi-player gaming system within each local path, and introduce Nash Chain-of-Thought (Nash CoT). Specifically, for a given question, we leverage LLM to autonomously select the contextually relevant template and generate outputs guided by this template, aiming to reach Nash Equilibrium alongside normal generation in each path. This approach allows us to achieve comparable or improved performance compared to self-consistency while using fewer inference paths on various inference tasks, including Arabic reasoning, Commonsense Question answering, and Symbolic inference.
@@
Ziqi Zhang1thanks:    Equal contribution.,  Cunxiang Wang [^1] footnotemark: 1,  Xiong Xiao2,  Yue Zhang1thanks:    Corresponding authors.,  Donglin Wang122footnotemark: 2 1 School of Engineering, Westlake University 2 University of Cambridge {zhangziqi, wangcunxiang, zhangyue, wangdonglin}@westlake.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) (Ouyang et al., 2022; etc., 2023; Jiang et al., 2023; Brown et al., 2020b; OpenAI, 2024). In particular, leveraging human-designed instructions as input, LLM demonstrates superior inference performance across various types of simple reasoning tasks (Radford et al., 2019; Brown et al., 2020a). But, its performance remains variable in complex reasoning tasks (Rae et al., 2022). To enhance LLM’s inference capabilities on complex issues, we can employ a step-by-step inference approach, known as Chain-of-Thought (CoT) prompting (Wei et al., 2023). For instance, starting with a template like "Let’s think step by step", LLM first generates rationales and then arrives at a final prediction. This approach significantly improves LLM’s inference performance across tasks like Arabic, Symbol Inference, and CommonsenseQA.
Figure 1: Demonstrations of Self-Consistency and Nash Chain-of-Thought (Nash CoT). (a) Self-Consistency entails inferring multiple paths and subsequently voting the most frequently prediction. (b) In Nash CoT, inference occurs multiple times, with each path generating two responses, but only one reached Preference Equilibrium is sustained. Ultimately, the answer sampled is the one that reaches preference equilibrium.
Subsequently, the impressive performance of CoT on complex inference tasks has spurred new developments in this direction (Wang et al., 2023; Wei et al., 2023; Jin and Lu, 2023; Zhang et al., 2022; Shi et al., 2022). Among these developments, self-consistency (Wang et al., 2023) emerges as a competitive CoT approach. It leverages uncertainty from multiple inference paths from a single LLM, and ranking generated answers by frequency can significantly enhance LLM’s inference performance. This approach significantly improves the performance of GPT-3 utilizing zero-shot CoT without any options for parameter tuning. Meanwhile, experimental results indicate that inference performance improves with the number of generated samples (Wang et al., 2023), suggesting that the potential of LLM’s inference capabilities has not yet to be fully exploited. Although self-consistency is straightforward to implement and requires no additional turning, it has a significant drawback: substantially higher inference costs compared to directly utilizing CoT.
On the other hand, the performance improvements resulting from self-consistency imply that relying solely on single-path inference cannot fully harness the inference capabilities of LLM, and multi-path inference encompasses the potential correct answers. Therefore, it’s necessary to maintain this strategy. Meanwhile, numerous practical applications indicate that when a LLM is given the appropriate templates, it can perform specific tasks more proficiently. Hence, an intuitive strategy to reduce the number of inference paths in self-consistency is to use templates to guide the LLM to correctly infer each path.
@@
To achieve this goal, we use the contextual information required by the question as a template to guide the inference in each path. This approach increases the probability that the LLM can correctly solve the question in each path, thereby potentially reducing the number of inference paths needed for multi-path inference. Meanwhile, to alleviate overconfidence in template-guided generation, we develop a bi-player gaming system that introduces the Nash Equilibrium of preference of strategy (defined as Preference Equilibrium) in multi-path inference. This system samples generations that balance the preferences of both template-guided and normal generation, ensuring the output robustly matches the context of the given question while sustaining some moderate generation pattern. Subsequently, we combine multi-path inference with Preference Equilibrium to propose Nash CoT. We present the comparison between Self-Consistency and Nash CoT in Figure 1.
We conduct experiments with two local deployed LLMs-Mistral-Instruct (Jiang et al., 2023) and GLM4 (Zeng et al., 2022; Du et al., 2022) on various inference tasks, including Arabic Reasoning (Koncel-Kedziorski et al., 2015; Hosseini et al., 2014) and Symbolic Reasoning (Wei et al., 2023) and Commonsense Reasoning (Talmor et al., 2019; Geva et al., 2021). As shown in Figure 2, Nash CoT has achieved similar or even better performance against self-consistency with fewer inference paths. Meanwhile, as shown in Figure 3, Nash CoT has significantly reduced the inference cost by up to 50% on local deployed LLMs.
To our knowledge, we are the first to introduce Preference Equilibrium into multi-path inference with the aim of harmonizing text generation guided by the optimal template with that generated by the model’s default state. It can reduce the number of paths needed to achieve good results in multi-path inference. And we are also the first to integrate Preference Equilibrium into multi-path inference and propose Nash CoT. It can achieve similar or even superior performance on various reasoning tasks while requiring only half costs compared to self-consistency. [^1] We will release our code at ¡https://github.com/stevezhangzA/nash-chain-of-thought¿.
@@ 2 Related Work
There are three majority kinds of CoT approaches. Zero-shot CoT that Prompts LLM with simple yet instructions to guide LLM generate step by step (Kojima et al., 2023). Manual CoT that initialized by randomly sampled several cases from dataset or designed by human, and followed by utilizing these cases as demonstration to guided LLM generate follow the manner of demonstration (Wei et al., 2023), however, such methods can be biased if the demonstration can represent real distribution. Automatic (or batch) CoT (Zhang et al., 2022) first sample cases which have the most representative sentence embedding in each clusters, followed by inference with the same manner as manual CoT. Self-Consistency (Wang et al., 2023) showcases strong performance in vast benchmarks. Apart from its impact on inference performance,self-consistency also boasts scalability as a key advantage. It seamlessly integrates with different approaches, such as tabular consistency of transformations (tab-CoT) Jin and Lu (2023), making it adaptable and versatile for various applications. Despite self-consistency has improved LLM’s performance on Arabic benchmarks. Correspondingly,self-consistency should have to inference multi-times, thus it burdens the deploy budgets.
One way to address this limitation is by initially sampling multiple inference paths and then fine-tuning using the highest frequency path. Specifically, Huang et al. (2022) suggest that gathering inferences from multiple paths and sampling the most frequent generation to fine-tune a smaller LLM can enhance the LLM’s inference performance. However, this approach still requires updating the LLM’s parameters, which is inefficient. Therefore, it is necessary to further explore inference methods to maintain the performance of self-consistency while reducing the number of inference paths.
@@
The training policy with Reinforcement Learning (RL) to reflect preference, termed Reinforcement Learning with Human Feedback (RLFH), was initially introduced by Akrour et al. (2011) and has since undergone consistent improvement and refinement by Cheng et al. (2011); Busa-Fekete et al. (2013); Wilson et al. (2012). This approach has been widely applied to adjust Large Language Models’ (LLMs) parameters to align with human preferences (Ouyang et al., 2022; Jiang et al., 2023; etc., 2023). Recently, a new approach called Direct Optimizing from Preference (DPO) has been proposed by Rafailov et al. (2023), aiming to directly adjust LLMs to reflect human preferences without requiring a reward model and RL algorithm. Additionally, Munos et al. (2023) proposed combining DPO with Nash equilibrium to ensure convergence of the last iterated policy. Our study also utilizes the concept of equilibrium in preference model, but the main difference is that we utilize preference equilibrium as a standard to pick up the most preferred answer instead optimizing the LLM’s parameters.
@@ 3 Preference Equilibrium in mini-batch inference
Self-consistency has shown that the inference of LLMs under a single path may not represent their full capabilities. By simply conducting multiple inferences and filtering answers through voting, it is possible to achieve more accurate results. However, multi-path inference lacks a strong theoretical foundation to determine the optimal number of inference paths, potentially leading to much more computational resource consumption. To reduce the number of paths required for multi-path inference, we utilize the concept of Nash Equilibrium to locally construct a binary game system in multi-path inference. Specifically, the preference of each valid inference path of the LLM needs to achieve Nash equilibrium with the preferences of the generation guided by the template. This approach increases the probability of each path correctly answering the question while maintaining a certain level of robustness, thereby reducing the number of inference paths required by self-consistency.
@@
Given text input $x$, and the sampled predictions (answers) $y_{1}$, $y_{2}$, we first define $y_{1}$ prefers $y_{2}$ as Equation 1:
$\begin{split}\mathcal{P}(y_{1}\prec y_{2}|x):={\rm sign}(r_{\theta}(y_{1}|x))-{\rm sign}(r_{\theta}(y_{2}|x)),\end{split}$ (1)
[…]
Player Templates: Templates for our preference model that are shown the structure: id(player): description. Mathematician: You are a mathematician, you excel at analyzing problems from a mathematical logical perspective and arrive at conclusions that align with your values. Literary scholar: You are a literary scholar who has read a vast array of literary works. Please consider the problem from the perspective of a literary scholar.
@@
Additionally, drawing from the definition from Munos et al., we define one policy as more preferred over another as:
$\begin{split}\mathcal{P}(\pi_{1}\prec\pi_{2}):=\mathbb{E}_{{}^{y_{1}\sim\pi_{1}(\cdot|x)}_{y_{2}\sim\pi_{2}(\cdot|x)}}[\mathcal{P}(y_{1}\prec y_{2}|x)]\end{split}$ (2)
@@
We aim to select the best template (we provide several cases in above example 3 for the current problem, thus facilitating the large model in problem-solving, correspondingly reduces the required num of inference paths. However, this may lead to some issues: 1) If the template is incorrectly chosen, it may cause the agent to generate answers outside the range of correct responses corresponding to the current problem, resulting in erroneous replies. 2) The large model may excessively generate context-dependent responses, affecting its robustness. To address these issues, we build a local bi-player gaming system that the preference of template guided LLM (player 1) over normal status of LLM (player 2) is the pay-off of template guided LLM, vice visa. If player 1 and player 2 reaches Nash Equilibrium  [^2] Preference Equilibrium leverages the concepts of Nash Equilibrium. Nash equilibrium is proposed by John Nash. It is a concept solution where, assuming each participant knows the equilibrium strategies of the other participants, no participant can benefit by changing their own strategy., then the strategy can match the preference of both player 1 and player 2.
Subsequently, we define the status that player 1 and player 2 reach Nash Equilibrium as Preference Equilibrium (Definition 1). Meanwhile, in Theorem 3.1, we prove the existence of Nash Equilibrium in this system. Specifically, the strategy of player 1 equal to player 2 is one solution that this system has reached Preference Equilibrium.
@@
Given any two policy (player) $\pi_{1}$ and $\pi_{2}$ within the gaming system defined in Definition 1, where $\pi\in\Pi$. $\pi_{1}\equiv\pi_{2}$denotes a solution where the gaming system reaches Nash Equilibrium.
Proof of Theorem 3.1 see Appendix E.
@@
Theorem 3.1 proves the existence of an Nash Equilibrium between the template guided LLM and the normal status of LLM. When reaching Preference Equilibrium, the preference of decisions made by the template guided LLM are aligned with those made by the LLM under normal status. Meanwhile, the preference of template guided LLM generation is much more closed to the requirement of quesitons’ context, while those of the normal status LLM are predominantly based on its parameters which is much more robust than template guided LLM. Therefore, this equilibrium can balance the requirement of contextual information and robustness of the model generation during problem-solving. Notabaly, $\pi_{1}\equiv\pi_{2}$ means their outputs are also likely to be equal. This insight forms a fundamental basis for piratically implementing Nash CoT.

### v2

@@ Abstract
Chain of thought (CoT) is a reasoning framework that can enhance the performance of Large Language Models (LLMs) on complex inference tasks. In particular, among various studies related to CoT, multi-path inference stands out as a simple yet effective improvement. However, there is no optimal setting for the number of inference paths. Therefore, we have to increase the number of inference paths to obtain better results, which in turn increases the inference cost. To address this limitation, we can utilize question-related role templates to guide LLMs into relevant roles, thereby increasing the possibility of correct inferences for each path and further reducing dependence on the number of inference paths while improving reasoning accuracy. However, placing LLMs into specific roles may reduce their reasoning diversity and performance on a few tasks where role dependence is low. To alleviate the excessive immersion of the LLM into a specific role, we propose Nash CoT by constructing a competitive system on each path that balances the generation from role-specific LLMs’ and the general LLMs’ generation, thereby ensuring both effective role adoption and diversity in LLM generation further maintaining the performance of multi-path inference while reducing the requirement of the number of inference paths. We evaluate Nash CoT across various inference tasks, including Arabic Reasoning, Commonsense Question Answering, and Symbolic Inference, achieving results that are comparable to or better than those of multi-path CoT with the equal number of inference paths.
@@
Ziqi Zhang1thanks:    Equal contribution.,  Cunxiang Wang [^1] footnotemark: 1,  Xiong Xiao2,  Yue Zhang1thanks:    Corresponding authors.,  Donglin Wang122footnotemark: 2 1 School of Engineering, Westlake University 2 University of Cambridge, UK stevezhangz@163.com {wangcunxiang, zhangyue, wangdonglin}@westlake.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have profoundly revolutionized the field of Natural Language Processing (NLP) (Ouyang et al., 2022; etc., 2023; Jiang et al., 2023; Brown et al., 2020b; OpenAI, 2024). Specifically, leveraging human-designed instructions as input, LLMs demonstrate superior inference performance across various types of simple reasoning tasks (Radford et al., 2019; Brown et al., 2020a).
However, in complex tasks, direct reasoning LLMs do not yield good results (Rae et al., 2022). To improve LLMs’ inference performance on complex inference tasks, it is popular to adopt a step-by-step reasoning framework known as Chain-of-Thought (CoT) prompting (Wei et al., 2023). For example, when appending templates that can guide LLMs to perform step-by-step reasoning, such as ‘‘Let’s think step by step", to the end of given question. We can lead the LLMs to output question-related rationale and further obtain the answer, enabling LLMs to achieve better performance than zero-shot reasoning.
Figure 1: Demonstrations of Nash Chain-of-Thought (Nash CoT). As shown in this figure, Nash CoT can be divided into three main steps. Step 1 involves bringing the LLM into a template-related role. Step 2 utilizes the role-immersed LLM and LLM under normal conditions to collect model predictions separately. Step 3 filters the responses based on Nash Equilibrium.
Subsequently, among various CoT-related improvements, modified by utilizing multi-path inference which is represented by Wang et al. (2023) is one of the most effective reasoning approaches. However, there is no theoretical proof identifying an optimal number of paths for multi-path reasoning. To ensure better inference performance, it is imperative to augment the inference paths. However, this might burden the inference budgets.
In order to reduce the dependency of multi-path inference on the number of inference paths while further enhancing its effectiveness, we can cast LLMs into a relevant role for a specific question. For instance, we can utilize a template such as "You’re a mathematician, and $\cdots$" to cast the LLM in the role of a mathematician. This approach can improve the accuracy of each path’s answer generated by the LLM, thereby reducing the multi-path CoT’s reliance on the number of paths, potentially improving the LLM’s reasoning.
However, excessively casting the LLM into a specific role using templates may compromise its robustness (diversity), thereby adversely affecting its performance in solving problems with minimal role dependency. To address this problem, we can introduce a certain degree of randomness into the reasoning process of the LLM. This can mitigate the constraints imposed by excessive role casting on the reasoning performance, ultimately maintaining the role-immersed multi-path CoT’s reasoning performance in these questions. Driven by this motivation, we construct a gaming system on each path, where the LLM in the role-playing context and the LLM in its normal state act as two distinct players. When this gaming system formed by these two players reaches a Nash Equilibrium, the reasoning under that path can balance the preferences of both role-playing with that of the LLM’s normal state. Furthermore, we can choose the response with the most Nash Equilibrium hits through voting to ensure the performance of LLM reasoning.
@@
We conduct experiments with two local deployed LLMs that include Mistral-Instruct (Jiang et al., 2023) and GLM-4 (Zeng et al., 2022; Du et al., 2022) on various inference tasks, including Arabic Reasoning (Koncel-Kedziorski et al., 2015; Hosseini et al., 2014), Symbolic Reasoning (Wei et al., 2023) and Commonsense Reasoning (Talmor et al., 2019; Geva et al., 2021). As shown in Figure 2, Nash CoT can achieve similar or even better performance against self-consistency with fewer inference paths. Meanwhile, as shown in Figure 3, Nash CoT can significantly reduce the inference cost by up to 50% on locally deployed LLMs. To summarize, our contribution can be summarized as follows:
  • To our knowledge, we are the first to introduce the concept of Nash Equilibrium into multi-path CoT, with the objective of balancing text generation guided by the role-immersed LLM with that produced by the LLM’s default state. This innovation enables us to minimize the number of paths required to achieve good results in multi-path inference, while simultaneously preserving the multi-path CoT’s inherent performance advantages.
  • We evaluated the performance of Nash CoT on a wide range of reasoning tasks. Nash CoT can achieve similar results to self-consistency in Arabic Reasoning, Symbolic Reasoning, and Commonsense Question Answering tasks, given the same number of reasoning paths. [^1] We have released our code at ¡https://github.com/stevezhangzA/nash-chain-of-thought¿.
@@ 2 Related Work
There are three majority types of CoT approaches. The first is zero-shot CoT, which involves prompting the LLM with simple instructions to guide its generation step by step (Kojima et al., 2023). The second type is Manual CoT, which begins with sampling several cases from a dataset or manually designing templates. These cases are then used as demonstrations to guide the LLM in generating responses (Wei et al., 2023). However, such methods can be biased if the demonstrations aren’t sufficient to cover the optimal distribution. Automatic CoT (Zhang et al., 2022) first clusters the dataset, then samples the most representative sentence from each cluster as the demonstrations and finally uses these demonstrations to guide LLM inference. On the other hand, self-consistency (Wang et al., 2023) showcases strong performance in vast benchmarks. Apart from its impact on inference performance, self-consistency also boasts scalability as a key advantage. It can seamlessly integrate with different approaches, such as tabular chain-of-thought (Tab-CoT) Jin and Lu (2023), making it adaptable and versatile for various applications. Furthermore, despite self-consistency can improve LLM’s performance on Arabic benchmarks, self-consistency must be inferred multiple times, burdening deployment budgets.
One way to overcome this limitation is to infer multiple paths and fine-tune them based on the most frequent path. Specifically, Huang et al. (2022) proposes that by gathering inferences from multiple paths and sampling the most frequent generation, we can enhance the inference performance of a smaller LLM. However, this approach still requires updating the LLM’s parameters, which is inefficient. Therefore, it is necessary to further develop inference methods to maintain the performance of self-consistency while reducing the number of multi-path CoT inference paths.
@@
Training policy with Reinforcement Learning (RL) to reflect preference, termed Reinforcement Learning with Human Feedback (RLFH), is initially introduced by Akrour et al. (2011) and has since undergone consistent improvement and refinement by Cheng et al. (2011); Busa-Fekete et al. (2013); Wilson et al. (2012). This approach has been widely applied to adjust the parameters of LLMs in order to align them with human preferences (Ouyang et al., 2022; Jiang et al., 2023; etc., 2023). Recently, a new approach called Direct Optimizing from Preference (DPO) has been proposed by Rafailov et al. (2023), aiming to directly adjust LLMs to reflect human preferences without requiring a reward model. Additionally, Munos et al. (2023) proposes combining DPO with Nash Equilibrium to ensure the convergence of the last iterated policy. Despite our study also utilizing the concept of Nash Equilibrium in the Preference model, the main difference is that we utilize Nash Equilibrium as a standard to pick up the answer instead of optimizing the LLM’s parameters.
@@ 3 Preference Equilibrium in mini-batch inference
The experimental results of self-consistency indicate that multi-path reasoning can unleash the reasoning capability of LLMs. However, multi-path inference lacks a theoretical foundation for determining the optimal number of inference paths, requiring us to increase the number of paths to ensure performance, and further increase the computational resource consumption. To reduce the number of paths required for multi-path inference, we utilize the concept of Nash Equilibrium to locally construct a binary gaming system in multi-path inference. Specifically, the preference of each valid inference path of the LLM needs to achieve Nash equilibrium with the preferences of the generation guided by the role-immersed LLM. This approach increases the possibility of each path correctly answering the question while maintaining a certain level of robustness (diversity), thereby reducing the number of inference paths required by self-consistency.
@@
Given text input $x$, and the sampled predictions (answers) $y_{1}$, $y_{2}$, and Bradley-Terry reward model $r_{\theta}$ that has been deliberated by Chen et al. We first define $y_{1}$ is preferred over $y_{2}$ as Equation 1:
$\begin{split}\mathcal{P}(y_{2}\prec y_{1}|x):={\sigma}(r_{\theta}(y_{1}|x)-r_{\theta}(y_{2}|x)),\end{split}$ (1)
where $\mathcal{P}$ is preference model reflects the preference when inputting the pairs $(y_{1},y_{2})$.
Player Templates : Role templates for our LLM that are shown the structure: {id(player): description.} Mathematician: You are a mathematician, you excel at analyzing problems from a mathematical logical perspective and arrive at conclusions that align with your values. Literary scholar: You are a literary scholar who has read a vast array of literary works. Please consider the problem from the perspective of a literary scholar.
@@
Furthermore, drawing from the definition in (Munos et al., 2023), we define one policy (LLM for decision making) $\pi_{1}$ as more preferred over another policy $\pi_{2}$ as:
$\begin{split}\mathcal{P}(\pi_{2}\prec\pi_{1}):=\mathbb{E}_{{}^{y_{1}\sim\pi_{1}(\cdot|x)}_{y_{2}\sim\pi_{2}(\cdot|x)}}\bigg{[}\mathcal{P}(y_{2}\prec y_{1}|x)\bigg{]},\end{split}$ (2)
subsequently, we imply the necessity and existence of Nash Equilibrium in the Preference model (shown in Equation 1).
@@
We first discuss why we should construct a gaming system: To increase the accuracy of each inference path, we can select the most suitable template (example 3) for the given question, immersing the LLM into a template-related role. This approach helps LLMs to solve problems more effectively, thereby reducing the number of inference paths needed.
However, this may lead to some issues: If the LLM is excessively immersed in a specific role, it may reduce the diversity of LLM’s generation, and further reduce LLM’s accuracy in solving tasks that have low role dependence. To address these problems, we locally build a bi-player gaming system that the preference of role-immersed LLM (player 1) over the normal status of LLM (player 2) is the pay-off of role-immersed LLM, vice visa. If player 1 and player 2 reach Nash Equilibrium  [^2] Preference Equilibrium leverages the concepts of Nash Equilibrium. Nash equilibrium is proposed by John Nash. It is a concept solution where, assuming each participant knows the equilibrium strategies of the other participants, no participant can benefit by changing their own strategy., then the generation can balance the preference of both player 1 and player 2.
Definition 3.1 (Preference Equilibrium).
Given any two policys (players) $\pi_{1}$ and $\pi_{2}$ along with the pay-off of $\pi_{1}$ and $\pi_{2}$ as $\mathcal{P}(\pi_{1}\prec\pi_{2})$ and $\mathcal{P}(\pi_{2}\prec\pi_{1})$, we define the Nash Equilibrium of $\pi_{1}$ and $\pi_{2}$ as Preference Equilibrium.
Subsequently, we define the status that player 1 and player 2 reach Nash Equilibrium as Preference Equilibrium (Definition 3.1). Meanwhile, in Theorem 3.1, we prove the existence of Nash Equilibrium in this system. In particular, the strategy of player 1 equal to player 2 is one solution that this system has reached Preference Equilibrium.
@@
Given any two policy (player) $\pi_{1}$ and $\pi_{2}$ within the gaming system defined in Definition 1, where $\pi\in\Pi$. $\pi_{1}\equiv\pi_{2}$ is a solution where the gaming system reaches Nash Equilibrium.
Proof of Theorem 3.1 see Appendix F.
@@
Theorem 3.1 proves the existence of a Nash Equilibrium between the role-immersed LLM and the normal status of LLM generation. Therefore, when reaching Preference Equilibrium, the preference of generations made by the role-immersed LLM is aligned with those made by the LLM without templates. Meanwhile, the preference for role-immersed LLM generation is much closer to the selected role-immersed context, while those of the normal status LLM are predominantly based on its parameters which are much more diverse than role-immersed LLM generation. Therefore, Preference Equilibrium can balance the requirement of role-immersed LLM generation and the diversity of the LLM generation in each path. In particular, $\pi_{1}\equiv\pi_{2}$ means their outputs are also likely to be equal. This insight forms a fundamental basis for practically implementing Nash CoT.